# SC-LoRA: Balancing Efficient Fine-tuning and Knowledge Preservation via Subspace-Constrained LoRA

### v2

@@ 6 Limitations
Although SC-LoRA can successfully handle both efficient fine-tuning and knowledge preservation at the same time, it still has drawbacks.
First, SC-LoRA is just a LoRA initialization method, and does not strongly constrain the updates during fine-tuning process. Hence after fine-tuning on more complex tasks and with more steps, the knowledge preservation ability can also drop (see the preservation drop of NQ-open in Table 3 for example).
Second, its application on preserving other types of knowledge remains unexplored. Future work may consider applying SC-LoRA to preserving multimodal large language model’s performance on pre-training tasks (Zhai et al., 2024) or large language model’s reasoning ability.
@@ Acknowledgments
Appendix A Definitions and Proofs
A.1 Mathematical Definition of $\Pi_{S}$
Definition 2.
$\Pi_{S}(x)$ $=\sum_{i=1}^{r}(q_{i}^{\top}x)q_{i}$ (10)
$=\sum_{i=1}^{r}(q_{i}q_{i}^{\top})x,\forall x\in\mathbb{R}^{n}.$
Note: the selection of the orthonormal basis does not affect $\Pi_{S}$.
A.2 Proof for Theorem 1
@@
From property of orthogonal matrix, $\sum_{i\in[r]}U_{ij}^{2}\leq 1$ and $\sum_{j=1}^{d}\sum_{i\in[r]}U_{ij}^{2}=r$, then to maximize $R$, from the additional assumption we need $\sum_{i\in[r]}U_{ij}^{2}=\begin{cases}1,&1\leq j\leq r\\ 0,&r+1\leq j\leq d\end{cases}$, which is equivalent to
@@ A.3 Proof of Theorem 2
Appendix B Numerical instability in sparse sample setting
@@
When the sample size is much larger than the output activation dimension, $\min(|\mathcal{D}_{+}|,|\mathcal{D}_{-}|)\gg d_{\mathrm{out}}$, setting $\beta\in[0,1]$ causes no issue. However, when samples are sparse (specifically, when the number of negative-task sample $N_{-}<d_{\mathrm{out}}-r$, setting $\beta=1$ introduces multiple valid solutions in the spectral decomposition step due to high-dimensional freedom in the null space of $\mathrm{Cov}_{-}$. Mathematically, the rank of $\mathrm{Cov}_{-}$ is at most $N_{-}$, resulting in a null space of dimension $d_{\mathrm{out}}-\mathrm{rank}(\mathrm{Cov}_{-})\geq d_{\mathrm{out}}-N_{-}>r$. Consequently, any arbitrary set of $r$ orthonormal vectors in this null space can satisfy the decomposition criterion, leading to non-unique initialization of parameters $A$ and $B$.
@@
To mitigate this instability, we recommend setting $1-\beta$ to a small positive value (rather than exactly zero). This retains the regularization from $\mathrm{Cov}_{+}$ in the objective function, which constrains the null space ambiguity and stabilizes the spectral decomposition, empirically improves fine-tuning performance.

### v3

@@ 4.4 Ablation: Experimental Analysis on the Functionality of Hyper-Parameter $\beta$
4.5 Ablation: Number of Samples for Initialization
A key aspect of our data-driven initialization is the number of samples used to estimate the covariance matrices. A robust method should not require an impractically large number of samples to be effective. In this section, we analyze the sensitivity of SC-LoRA’s performance to the number of initialization samples.
We conduct this analysis under two scenarios described earlier: data poisoning (Section 4.3) and benign fine-tuning (Section 4.2). For both scenarios, we vary the number of samples for both the fine-tuning task ($D_{+}$) and the preserved safety task ($D_{-}$) from $32$ to $256$, keeping all other hyperparameters constant ($\beta=0.9$, learning rate 2e52e-5). The results are presented in Table 5. From the table, we can draw several key conclusions: (1) Safety Preservation Improves with More Samples. This is expected, as a larger sample set provides a more accurate and stable estimate of the true covariance of the data distributions, allowing our method to better identify and constrain the safety-critical subspace. (2) Utility Remains Stable and Insensitive to Sample Size. For both scenarios, the utility score only fluctuates within a range of $1\%$. A possible reason is that even a small number of samples (like $32$) is sufficient to capture the principal directions of the fine-tuning task ($T_{+}$). Hence adding more samples may not significantly alter the subspace chosen for adaptation, leading to consistent performance on the downstream task.
While more samples are generally better for safety, the results suggest that for practical applications, using $128$ to $256$ samples provides a reliable balance between computational cost during initialization and robust performance. This is a practical and accessible number, reinforcing the efficiency of our approach.
@@ 6 Limitations
First, SC-LoRA is just a LoRA initialization method, and does not strongly constrain the updates during fine-tuning process. Hence after fine-tuning on more complex tasks and with more steps, the knowledge preservation ability can also drop (see the preservation drop of NQ-open in Table 1 for example). Second, its application on preserving other types of knowledge remains unexplored. Future work may consider applying SC-LoRA to preserving multimodal large language model’s performance on pre-training tasks (Zhai et al., 2024) or large language model’s reasoning ability.
@@ Acknowledgments
Appendix A Proof Details
Figure 4: Sensitivity to the rank $r$ on Samsum.
A.1 Proof for Theorem 1
@@
From the property of orthogonal matrix, $\sum_{i\in[r]}U_{ij}^{2}\leq 1$ and $\sum_{j=1}^{d}\sum_{i\in[r]}U_{ij}^{2}=r$, then to maximize $R$, from the additional assumption we need $\sum_{i\in[r]}U_{ij}^{2}=\begin{cases}1,&1\leq j\leq r\\ 0,&r+1\leq j\leq d\end{cases}$, which is equivalent to
@@ A.2 Proof of Theorem 2
Appendix B Ablation Study on Low Rank $r$
This section studies the sensitivity to the rank $r$ under experiment 4.2 (Benign Finetuning on Samsum). The results are plotted in Figure 4. Generally, SC-LoRA with $\beta=0.9$ achieves almost the best utility with small safety degradation on ranks from 16 to 128.
Appendix C Learning Rate Fine-tuning
This section presents the Pareto plots of learning rate fine-tuning and $\beta$ fine-tuning. See Figure 5 and Figure 6. The curve is closer to the bottom right, the safer response and better performance the model presents. From the Pareto plot, our SC-LoRA curve can best balance safety preservation and utility learning compared to other methods.
[…]
Data poisoning 256 1.136 1.212 45.26
128 1.155 2.121 45.64
64 1.200 2.727 46.10
32 1.864 1.616 45.72
Benign finetuning 256 1.097 0.000 51.67
128 1.118 1.515 51.91
64 1.145 1.515 52.46
32 1.158 1.515 52.42
Table 5: Sensitivity to the Number of Initialization Samples.
Appendix D Number of Samples in Initialization
This section compares the finetuning performance on different number of samples used to estimate the covariance matrices for SC-LoRA’s initialization. We compare under the experiment 4.3 (Safety Preservation on Data Poisoning Attack) and 4.2 (Safety Preservation on Benign Finetuning).
The results shows that performance remains relatively stable and effective for both tasks using 128 or more samples. Number of samples within 64 causes degradation in preservation.
Appendix E Ablation: Number of Samples for Initialization
A key aspect of our data-driven initialization is the number of samples used to estimate the covariance matrices. A robust method should not require an impractically large number of samples to be effective. In this section, we analyze the sensitivity of SC-LoRA’s performance to the number of initialization samples.
We conduct this analysis under two scenarios described earlier: data poisoning (Section 4.3) and benign fine-tuning (Section 4.2). For both scenarios, we vary the number of samples for both the fine-tuning task ($D_{+}$) and the preserved safety task ($D_{-}$) from $32$ to $256$, keeping all other hyperparameters constant ($\beta=0.9$). The results are presented in Table 5. From the table, we can draw several key conclusions: (1) Safety Preservation Improves with More Samples. This is expected, as a larger sample set provides a more accurate and stable estimate of the true covariance of the data distributions, allowing our method to better identify and constrain the safety-critical subspace. (2) Utility Remains Stable and Insensitive to Sample Size. For both scenarios, the utility score only fluctuates within a range of $1\%$. A possible reason is that even a small number of samples (like $32$) is sufficient to capture the principal directions of the fine-tuning task ($T_{+}$). Hence adding more samples may not significantly alter the subspace chosen for adaptation, leading to consistent performance on the downstream task.
While more samples are generally better for safety, the results suggest that for practical applications, using $128$ to $256$ samples provides a reliable balance between computational cost during initialization and robust performance. This is a practical and accessible number, reinforcing the efficiency of our approach.
Figure 5: Pareto plot of learning rate fine-tuning on task Samsum.
Figure 6: Pareto plot of learning rate fine-tuning on task poisoned math.
Appendix F Selection of hyperparameter $\beta$ in practice
@@
When the sample size of initialization is much larger than the output activation dimension, i.e. $\min(|\mathcal{D}_{+}|L,|\mathcal{D}_{-}|L)\gg d_{\mathrm{out}}$, setting $\beta\in[0,1]$ causes no issue. However when samples are sparse (specifically, when the number of negative-task sample $B_{-}<(d_{\mathrm{out}}-r)/L$, setting $\beta=1$ introduces multiple valid solutions in the spectral decomposition step due to high-dimensional freedom in the null space of $\mathrm{Cov}_{-}$. Mathematically, the rank of $\mathrm{Cov}_{-}$ is at most $B_{-}L$, resulting in a null space of dimension $d_{\mathrm{out}}-\mathrm{rank}(\mathrm{Cov}_{-})\geq d_{\mathrm{out}}-B_{-}L>r$. Consequently, any arbitrary set of $r$ orthonormal vectors in this null space can satisfy the decomposition criterion, leading to non-unique initialization of parameters $A$ and $B$.
In practice, $B_{-}\sim(d_{\mathrm{out}}-r)/L$, the decomposition results may also be affected by data selection and clipping, demonstrated by the drop of utility for $\beta=0.99$ (exhibited in Section 4.4).
@@
To mitigate this instability, we recommend setting $1-\beta$ to a small positive value (rather than exactly zero), for example $\beta=0.8$ or $0.9$ as discussed in Section 4. This retains the regularization from $\mathrm{Cov}_{+}$ in the objective function, which constrains the null space ambiguity and stabilizes the spectral decomposition, empirically improves fine-tuning performance.